OpenStudy (anonymous):

what is the inverse of 3 mod 5?

jimthompson5910 (jim_thompson5910):

We want to find the value of x that satisfies 3x = 1 (mod 5) 3*1 = 3 (mod 5) 3*2 = 6 = 1 (mod 5) So x = 2 (mod 5) Which means that the inverse of 3 mod 5 is 2.